[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun May 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun May 7 19:33:36 UTC 2017


HSU Shift Log for Sun May 07 2017 15:33:29 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: 2-day graphical TWO did not update on the web site.  
The 5-day outlook worked just fine.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170507/11d8785d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list